OpenStudy (ebonyserrano):

The residents of a city voted on whether to raise property taxes. The ratio of yes votes to no votes was 7 to 5. If there were 4210 no votes, what was the total number of votes?

OpenStudy (ebonyserrano):

I just need help with 2 more word problems

OpenStudy (will.h):

\[\frac{ 7 }{ 5 } = \frac{ 4210 - x }{ x }\] Cross multiply and notice that ratio 7:5 means 12/5 so therefor 12x = 5 * 4210 12x = 21050 x=1754.17 people who said no

OpenStudy (ebonyserrano):

it was actually 10,104 votes

OpenStudy (will.h):

that was the number who voted no add that to those who voted yes and you'll get the total votes

OpenStudy (brainlesslonelygirl):

 cross multiply and solve for x. that will be the number of yes votes. then add that to the number of no votes to get the total number of votes.